### v1

Abstract.
In [EV] we introduced the notion of a periodic pencil of flat connections on a smooth variety $X$. Namely, a pencil is a linear family of flat connections $\nabla(s_{1},...,s_{n})=d-\sum_{i=1}^{r}\sum_{j=1}^{n}s_{j}B_{ij}dx_{i}$, where $\{x_{i}\}$ are local coordinates on $X$ and $B_{ij}:X\to{\rm Mat}_{N}$ are matrix-valued regular functions. A pencil is periodic if it is generically invariant under the shifts $s_{j}\mapsto s_{j}+1$ up to isomorphism. In this paper we show that in characteristic $p$, the $p$-curvature operators $\{C_{i},1\leq i\leq r\}$ of a periodic pencil $\nabla$ are isospectral to the commuting endomorphisms $C_{i}^{*}:=\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{ij}^{(1)}$, where $B_{ij}^{(1)}$ is the Frobenius twist of $B_{ij}$. Using the results of [EV], this allows us to compute the eigenvalues of the $p$-curvature for many important examples of pencils of flat connections, including Knizhnik-Zamolodchikov, Dunkl, and Casimir connections, as well as equivariant quantum connections for conical symplectic resolutions with finitely many torus fixed points.
To Boris Feigin on his 70-th birthday with admiration
1. Introduction
Grothendieck’s $p$-curvature is an important invariant of linear ODE on algebraic curves and, more generally, of holonomic systems of linear PDE on algebraic varieties. For example, if such a system over $\mathbb{C}$ has an algebraic fundamental solution then its reduction at almost all primes $p$ has zero $p$-curvature, and the converse is the famous Grothendieck-Katz conjecture. Further, if such an irreducible system is geometric (i.e., occurs as a composition factor in a Gauss-Manin connection on the cohomology of fibers of a smooth morphism) then by a theorem of N. Katz its reduction at almost all primes $p$ has nilpotent $p$-curvature, and the converse is the André-Bombieri-Dwork conjecture.
However, for a general holonomic system the $p$-curvature (and even its spectrum) is hard to compute, even if $X$ is an open subset of $\mathbb{A}^{1}$. The goal of this paper is to show that nevertheless for many important holonomic systems the $p$-curvature can be studied effectively, and in particular one can compute its spectrum.
Namely, in [EV] we introduced the notion of a periodic pencil of flat connections and showed that many important connections, such as Knizhnik-Zamolodchikov (KZ), Dunkl, and Casimir connections, as well as equivariant quantum connections of conical symplectic resolutions with finitely many torus fixed points fall into this category. Our main motivation in introducing this notion was the computation of the spectrum of the $p$-curvature of such connections in positive characteristic, which is what we do in this paper. Namely, our main result (Theorem 3.3) states that the $p$-curvature operators $\{C_{i},1\leq i\leq r\}$ of a periodic pencil $\nabla(\mathbb{s})=d-\sum_{i=1}^{r}\sum_{j=1}^{n}s_{j}B_{ij}dx_{i}$ are isospectral to the commuting endomorphisms $\sum_{i=1}^{r}\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{ij}^{(1)}dx_{i}$, where $x_{i}$ are local coordinates on $X$, $B_{ij}:X\to{\rm Mat}_{N}$ are regular functions, and $B_{ij}^{(1)}$ is the Frobenius twist of $B_{ij}$. In particular, it follows that periodic pencils over $\overline{\mathbb{Q}}$ are globally nilpotent for $\mathbb{s}\in\mathbb{Q}^{n}$. Using this theorem, we compute the spectrum of the $p$-curvature for the above examples of periodic pencils and show (using a theorem of Katz) that every periodic pencil over $\mathbb{C}$ has regular singularities.
Theorem 3.3 also has an application to symplectic geometry. Namely, it is shown in [Lee] that if $X$ is a conical symplectic resolution with finitely many fixed points under the action of a torus $\mathbb{T}$, then the equivariant version of Fukaya’s quantum Steenrod operation ${\rm St}$ on $H^{*}_{\mathbb{T}}(X,\mathbb{F}_{p})$ coincides with the $p$-curvature $C$ of the equivariant quantum connection $\nabla(\mathbb{s})$ for $X$, provided the latter has simple spectrum (and conjecturally always). In any case, it is proved in [Lee] that these two operators commute and coincide after raising to power $p^{i}$ for some $i$, so in particular they are isospectral. On the other hand, the pencil $\nabla(\mathbb{s})$ is periodic due to existence of geometric shift operators (see [BMO, MO] and references therein), so our result yields the spectrum of $C$, and thereby the spectrum of ${\rm St}$.
The organization of the paper is as follows.
In Section 2 we discuss preliminaries.
In Section 3 we prove Theorem 3.3. We then apply this theorem to show that periodic pencils are globally nilpotent at rational parameter values and have regular singularities. Further, we introduce infinitesimally-periodic and mixed-periodic pencils, which arise as confluent limits of periodic pencils, and generalize Theorem 3.3 to such pencils. Next, we describe applications of Theorem 3.3 to periodic pencils from [EV]. Finally, we discuss the connection of our results with Katz’s theorem on global nilpotence of Gauss-Manin connections and with the André-Bombieri-Dwork conjecture.
Finally, in Section 4 we discuss irregular pencils and compute the spectrum of their $p$-curvature in several cases.
Acknowledgements. The authors thank Vadim Vologodsky and Jae Hee Lee for useful discussions. P. E.’s work was partially supported by the NSF grant DMS-2001318 and A. V.’s work was partially supported by the NSF grant DMS-1954266.
2. Preliminaries
Throughout the paper, $\mathbb{k}$ denotes an algebraically closed field, $V$ a finite dimensional $\mathbb{k}$-vector space, and $X$ a smooth irreducible algebraic variety over $\mathbb{k}$, unless specified otherwise.
2.1. Reduction to characteristic $p$
Let $Y$ be any algebro-geometric structure defined over $\overline{\mathbb{Q}}$ which depends on finitely many parameters (e.g., variety, morphism of varieties, vector bundle, connection, etc.). Then $Y$ can be defined over a finitely generated subring $R\subset\overline{\mathbb{Q}}$. More precisely, there exists a form $Y_{R}$ of $Y$ defined over $R$ such that $Y=Y_{R}\otimes_{R}\overline{\mathbb{Q}}$. For every prime $p$, there are finitely many homomorphisms $\phi:R\to\overline{\mathbb{F}}_{p}$ (at least one for almost all $p$). Given such $\phi$, we can define the corresponding reduction of $Y$ at $p$, ${\rm Red}_{p}(R,Y_{R},\phi):=Y_{R}\otimes_{R,\phi}\overline{\mathbb{F}}_{p}$. Of course, two different choices of $(R,Y_{R},\phi)$ may give isomorphic reductions. In fact, it is easy to show that for any two choices $(R^{\prime},Y_{R^{\prime}})$ and $(R^{\prime\prime},Y_{R^{\prime\prime}})$ the collection of reductions corresponding to various $\phi$ is the same for almost all $p$. Thus it makes sense to say that reductions of $Y$ at almost all $p$ enjoy a certain property.
2.2. Frobenius twists and morphisms
Let ${\rm char}(\mathbb{k})=p>0$. Define the Frobenius twist functor $V\mapsto V^{(1)}$ on the category ${\rm Vect}_{\mathbb{k}}$ of vector spaces over $\mathbb{k}$ by setting $V^{(1)}$ to be the subspace of $S^{p}V$ of elements $v^{p}$, $v\in V$ (this is an additive symmetric monoidal functor). It is clear that if $\{v_{i}\}$ is a basis of $V$ then $\{v_{i}^{p}\}$ is a basis of $V^{(1)}$.
For a linear map $T:V\to W$ let $T^{(1)}:V^{(1)}\to W^{(1)}$ be the image of $T$ under the Frobenius twist functor. Then, if $(t_{ij})$ is the matrix of $T$ in bases $\{v_{i}\}$, $\{w_{j}\}$ then the matrix of $T^{(1)}$ in the bases $\{v_{i}^{p}\}$, $\{w_{j}^{p}\}$ is $(t_{ij}^{p})$. Thus if $V=W$ are finite dimensional then the eigenvalues of $T^{(1)}$ are $\lambda_{j}^{p}$, where $\lambda_{j}$ are the eigenvalues of $T$.
We have a natural twisted-linear isomorphism $V\cong V^{(1)}$ which sends $v\in V$ to $v^{p}$. Thus we may alternatively define $V^{(1)}$ as the $\mathbb{F}_{p}$-space $V$ with twisted scalar multiplication given by $\lambda\cdot v:=\lambda^{1/p}v$. In this realization, $T^{(1)}$ is identified with $T$.
Now let $X$ be an affine scheme over $\mathbb{F}_{p}$. The absolute Frobenius of $X$ is the morphism ${\bf Fr}:X\to X$ induced by the $p$-th power homomorphism $\mathbb{F}_{p}[X]\to\mathbb{F}_{p}[X]$, $f\mapsto f^{p}$.
Suppose now that $X$ is an affine $\mathbb{k}$-scheme. Then the absolute Frobenius ${\bf Fr}:X\to X$ is not a morphism of $\mathbb{k}$-schemes, since the corresponding map on regular functions is not $\mathbb{k}$-linear but rather twisted $\mathbb{k}$-linear. To correct this, define the Frobenius twist $X^{(1)}$ of $X$ to be the affine $\mathbb{k}$-scheme with $\mathbb{k}[X^{(1)}]:=\mathbb{k}[X]^{(1)}$. Thus $X^{(1)}=X$ as an $\mathbb{F}_{p}$-scheme but with $\mathbb{k}$-structure twisted by the map $\lambda\mapsto\lambda^{1/p}$. This is a non-linear version of the Frobenius twist of vector spaces defined above.
Now the absolute Frobenius gives rise to a morphism of $\mathbb{k}$-schemes ${\rm Fr}:X\to X^{(1)}$ called the relative Frobenius of $X$, which on functions is defined by the same formula $f\mapsto f^{p}$. Note that ${\rm Fr}$ is a homeomorphism in Zariski topology, so we may identify $X$ with $X^{(1)}$ as topological spaces using ${\rm Fr}$.
By considering affine open covers, these definitions extend straightforwardly to not necessarily affine schemes.
2.3. Isospectrality
Let $R$ be a finitely generated commutative $\mathbb{k}$-algebra. Let $V$ be an $R$-module which is finite dimensional over $\mathbb{k}$. Then $V$ has a finite composition series consisting of simple $R$-modules, which must be of the form $R/\mathfrak{m}\cong\mathbb{k}$, where $\mathfrak{m}\subset R$ is a maximal ideal. Thus we obtain a function $\mu_{V}:{\rm Specm}R\to\mathbb{Z}_{\geq 0}$ with finite support such that $\mu_{V}(\mathfrak{m})$ is the multiplicity of $R/\mathfrak{m}$ in the Jordan-Hölder series of $V$. This function is called the spectral multiplicity of $V$, and its support is called the spectrum of $V$, denoted $\Sigma(V)$. If $\mu_{V}|_{\Sigma(V)}=1$, we say that $V$ has simple spectrum. It is easy to see that if $V$ has simple spectrum then it is semisimple (a direct sum of simple modules).
Two finite dimensional $R$-modules $V,W$ having the same spectral multiplicity are called isospectral. Let $V_{s}$ be the semisimplification of $V$, i.e., the direct sum of all simple composition factors of $V$. Thus $V,W$ are isospectral iff $V_{s}\cong W_{s}$.
Now let $U$ be a finite dimensional $\mathbb{k}$-vector space and $R:={\rm Sym}U$. Then an $R$-module is just a $\mathbb{k}$-vector space $V$ equipped with a linear map $L:U\to{\rm End}V$ such that
$[L(\mathbb{u}_{1}),L(\mathbb{u}_{2})]=0$
for all $\mathbb{u}_{1},\mathbb{u}_{2}\in U$. Thus if $\mathbb{u}=(u_{1},...,u_{r})$ in some basis of $U$ then $L(\mathbb{u})=\sum_{i=1}^{r}u_{i}L_{i}$, where $L_{1},...,L_{r}\in{\rm End}V$ are commuting linear operators. We call such $L$ a pencil of commuting endomorphisms, and say that $V$ corresponds to $L$, writing $V=V_{L}$. We say that two such pencils $L,M$ are isospectral if so are the modules $V_{L}$ and $V_{M}$.
Lemma 2.1.
$L,M$ are isospectral if and only if for every $\mathbb{u}\in U$, the characteristic polynomials of the operators $L(\mathbb{u})$ and $M(\mathbb{u})$ (or, equivalently, their eigenvalues counting multiplicities) are the same.
Proof.
The “only if” part is obvious, so let us prove the “if” part. Pick $\mathbb{u}\in U$ which separates points of $\Sigma(V_{L})\cup\Sigma(V_{M})\subset U^{*}$. This is possible, since any element of $R=\mathbb{k}[u_{1},...,u_{r}]$ vanishing on $\mathbb{k}^{r}$ is zero. For any $\ell\in\Sigma(V_{L})\cup\Sigma(V_{M})$, $\mu_{V_{L}}(\ell)$, respectively $\mu_{V_{M}}(\ell)$, equals the multiplicity of the eigenvalue $\ell(\mathbb{u})$ for the operator $L(\mathbb{u})$, respectively $M(\mathbb{u})$. Thus for such $\ell$, we have $\mu_{V_{L}}(\ell)=\mu_{V_{M}}(\ell)$, hence $\mu_{V_{L}}=\mu_{V_{M}}$, as desired. ∎
Example 2.2.
If ${\rm char}(\mathbb{k})=p$ and $L:=\sum_{i=1}^{r}u_{i}L_{i}:U\to{\rm End}V$ is a pencil of commuting endomorphisms, then the pencil $L(\mathbb{u})^{p}=\sum_{i=1}^{r}u_{i}^{(1)}L_{i}^{p}\in{\rm End}V$ parametrized by $U^{(1)}$ (where $u_{i}^{(1)}:=u_{i}^{p}$) is isospectral to the pencil $L(\mathbb{u})^{(1)}=\sum_{i=1}^{r}u_{i}^{(1)}L_{i}^{(1)}\in{\rm End}V^{(1)}$. Indeed, the eigenvalues of both are $\lambda_{j}(\mathbb{u})^{p}$, where $\lambda_{j}(\mathbb{u})$ are the eigenvalues of $L(\mathbb{u})$.
2.4. $p$-curvature
Let $\nabla$ be a flat connection on a vector bundle $\mathcal{V}$ over $X$. If $x_{1},...,x_{r}$ are local coordinates on $X$ then we have commuting operators $\nabla_{1},...,\nabla_{r}$ of covariant partial derivatives with respect to these coordinates, which act on the space $\Gamma_{\rm rat}(\mathcal{V})$ of rational sections of $\mathcal{V}$ and determine $\nabla$.
Suppose that ${\rm char}(\mathbb{k})=p>0$. Then the operators
$C_{i}=C_{i}(\nabla):=\nabla_{i}^{p}:\Gamma_{\rm rat}(\mathcal{V})\to\Gamma_{\rm rat}(\mathcal{V})$
are $\mathbb{k}(X)$-linear and combine into a twisted 1-form
$C=C(\nabla):=\sum_{i=1}^{r}C_{i}(\nabla)dx_{i}^{(1)}\in{\rm Fr}^{*}\Omega^{1}(X^{(1)})\otimes_{\mathcal{O}_{X}}{\rm End}\mathcal{V},$
where ${\rm Fr}:X\to X^{(1)}$ is the relative Frobenius morphism and $x_{i}^{(1)}:=x_{i}^{p}$ are local coordinates on $X^{(1)}$. The form $C$ is independent on the choice of coordinates and is called the $p$-curvature of $\nabla$. Moreover, $[\nabla,C]=[C,C]=0$, i.e., $[\nabla_{i},C_{l}]=[C_{i},C_{l}]=0$ for all $i,l$.
Example 2.3.
If $X=\mathbb{A}^{1}$ and $\nabla=d+a$, $a\in{\rm Mat}_{N}(\mathbb{k}[x])$, then for $p=2$ we have $C=a^{2}+a^{\prime}$ and for $p=3$ we have $C=a^{3}+[a^{\prime},a]+a^{\prime\prime}$ (where we identify 1-forms with functions using the coordinate $x$).
An important application of $p$-curvature is the computation of the space of local flat sections of $\nabla$ (see e.g. [K1]). Let us recall this computation. For $\mathbb{x}\in X(\mathbb{k})$ let $\mathcal{K}_{\mathbb{x}}\subset\mathcal{V}_{\mathbb{x}}$ be the kernel of the $p$-curvature $C(\nabla)(\mathbb{x})$. Assume that $\dim\mathcal{K}_{\mathbb{x}}=d$ for all $\mathbb{x}$ (this can be achieved by replacing $X$ by a dense open subvariety). Thus we have a vector bundle $\mathcal{K}$ on $X$ whose fibers are $\mathcal{K}_{\mathbb{x}}$. Let $\mathcal{S}$ be the sheaf of flat sections of $\nabla$.
Proposition 2.4.
The sheaf $\mathcal{S}$ is a locally free coherent sheaf of rank $d$ over $X^{(1)}$, and $\mathcal{K}={\rm Fr}^{*}\mathcal{S}$.
Proof.
Since $\nabla$ commutes with $C$, it defines a flat connection on $\mathcal{K}$. Moreover, if $F$ is a flat section of $\nabla$, i.e., $\nabla_{i}F=0$ for all $i$, then $C_{i}F=\nabla_{i}^{p}F=0$, so $F$ is a section of $\mathcal{K}$. Thus, restricting attention to $\mathcal{K}$, we may assume without loss of generality that $C=0$.
Let $\mathbb{x}\in X(\mathbb{k})$. On some affine neighborhood $U$ of $\mathbb{x}$ we can choose local coordinates $x_{1},...,x_{r}$ near $\mathbb{x}$ and identify $\mathcal{V}$ with the trivial bundle $\mathbb{k}^{N}\times U\to U$. Our job is to show that for any $F_{0}\in\mathbb{k}^{N}$ there exists a flat section $F(x_{1},...,x_{r})$ of $\nabla$ such that $F(0)=F_{0}$. But one such section is given by the formula
$F=(-1)^{r}\prod_{i=1}^{r}\nabla_{i}^{p-1}(F_{0}\prod_{i=1}^{r}x_{i}^{p-1}),$
which completes the proof. ∎
More concretely, Proposition 2.4 implies that if $\mathcal{K}_{\mathbb{y}}$ has constant dimension $d$ for $\mathbb{y}$ in a neighborhood of $\mathbb{x}\in X(\mathbb{k})$ then there exists an affine neighborhood $U$ of $\mathbb{x}$ and flat sections $F_{1},...,F_{d}$ of $\nabla$ over $U$ such that $F_{1}(\mathbb{y}),...,F_{d}(\mathbb{y})$ is a basis of $\mathcal{K}_{\mathbb{y}}$ for all $\mathbb{y}\in U(\mathbb{k})$, and $F_{1},...,F_{d}$ is a basis of $\mathcal{S}(U)$ over $\mathcal{O}(U^{(1)})$.
Remark 2.5.
Suppose $\nabla$ is a rational connection on $\mathbb{A}^{1}$ defined over $\overline{\mathbb{Q}}$. If $\nabla$ has a basis of algebraic solutions (which for regular connections is equivalent to having finite monodromy), then by Proposition 2.4 the $p$-curvature of the reduction of $\nabla$ to characteristic $p$ is zero for almost all $p$, since the algebraic solutions can be reduced modulo large $p$. The converse statement is known as the Grothendieck-Katz conjecture, which first appeared in [K1] and is still open.
3. Periodic pencils of flat connections and their $p$-curvature
3.1. Periodic families and pencils of flat connections
In [EV] we defined periodic families of flat connections. Let us recall this definition.
Definition 3.1.
An $n$-parameter family of flat connections on $X$ with values in $V$ is a family of flat connections $\nabla(\mathbb{s})=d-B(\mathbb{s})$, $\mathbb{s}:=(s_{1},...,s_{n})$, on the trivial vector bundle [^1] As explained in [EV], our theory can be generalized to arbitrary vector bundles on $X$. However, for simplicity we consider only trivial bundles. This assumption is harmless because in all applications the relevant bundle turns out to be trivial, and in any case it can always be trivialized by replacing $X$ with a dense open subset. $X\times V\to X$, where $B\in\Omega^{1}(X)\otimes{\rm EndV}[\mathbb{s}]$. The family $\nabla$ is said to be a pencil if $B=\sum_{j=1}^{n}s_{j}B_{j}$, $B_{j}\in\Omega^{1}(X)\otimes{\rm End}V$.
[…]
Definition 3.2.
A family $\nabla$ is said to be periodic if there exist shift operators
[…]
such that
[…]
Suppose ${\rm char}(\mathbb{k})=0$ and $\nabla$ is a periodic family over $\mathbb{k}$. In this case we can choose forms of $X$, $\nabla$ and $A$ over some finitely generated subring $S\subset\mathbb{k}$. Then for any homomorphism $\phi:S\to\mathbb{F}$ from $S$ to a field $\mathbb{F}$ of characteristic $p$, the family $\nabla\otimes_{S}\mathbb{F}$ obtained from $\nabla$ by reduction to $\mathbb{F}$ via $\phi$ is a periodic family over $\mathbb{F}$. Thus every example of a periodic family in characteristic zero gives a periodic family in almost every positive characteristic.
3.2. $p$-curvature of periodic pencils
Now assume that char$(\mathbb{k})=p$. Let $U_{X}$ be the vector space over $\mathbb{k}(X)$ of rational vector fields on $X$. Let $\nabla(\mathbb{s})=d-\sum_{j=1}^{n}s_{j}B_{j}$ be a pencil of flat connections on $X$. Then for each $\mathbb{s}=(s_{1},...,s_{n})\in\mathbb{k}^{n}$ we have two pencils of commuting endomorphisms: the $p$-curvature
$C(\nabla(\mathbb{s})):U_{X^{(1)}}\to{\rm End}V$
and
$\mathbb{B}(\nabla(\mathbb{s})):=\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{j}^{(1)}=\sum_{i=1}^{r}\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{ij}^{(1)}dx_{i}^{(1)}:U_{X^{(1)}}\to{\rm End}V^{(1)}.$
Our main result is the following theorem.
Theorem 3.3.
If $\nabla(\mathbb{s})$ is a periodic pencil then the pencils $C(\nabla(\mathbb{s}))$ and $\mathbb{B}(\nabla(\mathbb{s}))$ of commuting endomorphisms are isospectral for all $\mathbb{s}\in\mathbb{k}^{n}$. In particular, if $s_{j}\in\mathbb{F}_{p}$ for all $j$ then the endomorphisms $C_{i}(\nabla(\mathbb{s}))$ are nilpotent. Moreover, if for some $\mathbb{s}$, $\mathbb{B}(\nabla(\mathbb{s}))$ has simple spectrum (in the sense that $\mathbb{B}(\nabla(\mathbb{s}))(\mathbb{x})$ has simple spectrum for generic $\mathbb{x}\in X$), then $C(\nabla(\mathbb{s}))$ and $\mathbb{B}(\nabla(\mathbb{s}))$ are conjugate.
Proof.
For $u_{1},...,u_{r}\in\mathbb{k}$ let
$C(\mathbb{u},\mathbb{s}):=\sum_{i=1}^{r}u_{i}C_{i}(\nabla(\mathbb{s})),\ \mathbb{B}(\mathbb{u},\mathbb{s}):=\sum_{i=1}^{r}\sum_{j=1}^{n}u_{i}(s_{j}-s_{j}^{p})B_{ij}^{(1)}.$
It suffices to show that for all $\mathbb{u},\mathbb{s}$ the characteristic polynomials of the matrices $C(\mathbb{u},\mathbb{s})$ and $\mathbb{B}(\mathbb{u},\mathbb{s})$ coincide.
Fix $\mathbb{u}$ and let
$b_{m}(\mathbb{u},\mathbb{s}):={\rm Tr}\wedge^{m}C(\mathbb{u},\mathbb{s})$
be the $m$-th coefficient of the characteristic polynomial of $C(\mathbb{u},\mathbb{s})$. This is a polynomial of $\mathbb{s}$ of degree $pm$. The periodicity property of $\nabla$ implies that
(3.1) $C(\nabla(\mathbb{s}+\mathbb{e}_{j}))\circ A_{j}(\mathbb{s})=A_{j}(\mathbb{s})\circ C(\nabla(\mathbb{s})),$
hence $b_{m}(\mathbb{u},\mathbb{s}+\mathbb{e}_{j})=b_{m}(\mathbb{u},\mathbb{s})$ for all $j$. Thus
$b_{m}(\mathbb{u},\mathbb{s})=\beta_{m}(\mathbb{u},\mathbb{s}-\mathbb{s}^{p}),$
where $\mathbb{s}^{p}:=(s_{1}^{p},...,s_{n}^{p})$ and $\beta_{m}$ is a polynomial of degree $m$ in the second variable.
Moreover, it is easy to see that the leading term of $b_{m}(\mathbb{u},\mathbb{s})$ in $\mathbb{s}$ is
$b_{m}^{0}(\mathbb{u},\mathbb{s})={\rm Tr}\wedge^{m}\sum_{i=1}^{r}u_{i}(\sum_{j=1}^{n}s_{j}B_{ij})^{p}.$
In view of Example 2.2, we then have
$b_{m}^{0}(\mathbb{u},\mathbb{s})={\rm Tr}\wedge^{m}\sum_{i=1}^{r}u_{i}\sum_{j=1}^{n}s_{j}^{p}B_{ij}^{(1)}.$
Hence the leading term $\beta_{m}^{0}$ of $\beta_{m}$ is
$\beta_{m}^{0}(\mathbb{u},\mathbb{s})={\rm Tr}\wedge^{m}\sum_{i=1}^{r}\sum_{j=1}^{n}u_{i}s_{j}B_{ij}^{(1)},$
so
$\beta_{m}^{0}(\mathbb{u},\mathbb{s}-\mathbb{s}^{p})={\rm Tr}\wedge^{m}\mathbb{B}(\mathbb{u},\mathbb{s}).$
Thus it remains to show that $\beta_{m}=\beta_{m}^{0}$, i.e., $\beta_{m}$ is homogeneous of degree $m$ in $\mathbb{s}$. To this end it suffices to show that the polynomial $\beta_{m}(\mathbb{u},\mathbb{s}-\mathbb{s}^{p})$ does not contain monomials of degree $<m$ in $\mathbb{s}$. But $C(\mathbb{u},0)=0$, so all matrix coefficients of $C(\mathbb{u},\mathbb{s})$ belong to the ideal $(s_{1},...,s_{n})$. This implies that ${\rm Tr}\wedge^{m}C(\mathbb{u},\mathbb{s})$ does not contain monomials of degree $<m$, as desired. ∎
Thus the simultaneous eigenvalues $\Lambda_{i\ell}$ of the operators $C_{i}(\nabla(\mathbb{s}))$ coincide with the simultaneous eigenvalues of the operators $\mathbb{B}_{i}(\mathbb{s}):=\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{ij}^{(1)}$. If $n=1$ and $s_{1}=s$ then we have $\Lambda_{i\ell}=(s-s^{p})\Lambda_{i\ell}^{0}$, where $\Lambda_{i\ell}^{0}$ are the simultaneous eigenvalues of the operators $B_{i}^{(1)}$ which don’t depend on $s$. However, for $n\geq 2$ these eigenvalues are more complicated algebraic (in general, irrational) functions of $\mathbb{s}$, since $B_{ij}$ for different $j$ in general don’t commute.
Remark 3.4.
If $\mathbb{s}\in\mathbb{F}_{p}^{n}$ then $\mathbb{B}(\nabla(s))=0$ but the nilpotent 1-forms $C(\nabla(\mathbb{s}))$ need not be zero (except, of course, for $\mathbb{s}=0$). However, many of these forms are often conjugate to each other. For example, consider the case $n=1$, and let $a_{1},...,a_{\ell}\in\mathbb{F}_{p}$ be the poles of $A,A^{-1}$ which belong to $\mathbb{F}_{p}$, arranged in “increasing order”, i.e., when $a_{j}$ are regarded as elements of $[0,p-1]$, we have $10\leq a_{1}<a_{2}<...<a_{\ell}\leq p-$ 1. Then $\mathbb{F}_{p}$ is divided into intervals $I_{j}:=[a_{j}+1,a_{j+1}]$, $j\in\mathbb{Z}/\ell$ (here the interval $I_{0}$ consists of such $m\in[0,p-1]$ that $a_{\ell}+1\leq m\leq p-1$ or $0\leq m\leq a_{1}$). Equation (3.1) now implies
Proposition 3.5.
For every $j\in\mathbb{Z}/\ell$, the forms $C(\nabla(m))$, $m\in I_{j}$ are conjugate to each other. In particular, if $m\in I_{0}$ then $C(\nabla(m))=C(\nabla(0))=0$.
3.3. Infinitesimally periodic and mixed-periodic families
By replacing $\mathbb{s}$ with $\mathbb{s}/t$ in the definition of a periodic family of flat connections and sending $t$ to 00, we obtain the definition of an infinitesimally periodic family.
Definition 3.6.
A family of connections $\nabla(\mathbb{s})=d-B(\mathbb{s})$ on the trivial bundle on $X$ with fiber $V$ is infinitesimally periodic if there exist infinitesimal shift operators
[…]
such that
[…]
i.e., ~$\widetilde{\nabla}:=d-B(\mathbb{s})-\sum_{j=1}^{n}a_{j}(\mathbb{s})ds_{j}$ is a connection on the trivial bundle on $X\times\mathbb{k}^{n}$ with fiber $V$ whose curvature is of the form $\eta=\sum_{1\leq i<j\leq n}\eta_{ij}(\mathbb{x},\mathbb{s})ds_{i}\wedge ds_{j}$.
More generally, one may consider mixed-periodic families $\nabla(\mathbb{s})$, $\mathbb{s}=(\mathbb{s}_{1},\mathbb{s}_{2})$, $\mathbb{s}_{1}\in\mathbb{k}^{n_{1}}$, $\mathbb{s}_{2}\in\mathbb{k}^{n_{2}}$, $n_{1}+n_{2}=n$, which are periodic in $\mathbb{s}_{1}$ for fixed $\mathbb{s}_{2}$ and infinitesimally periodic in $\mathbb{s}_{2}$ for fixed $\mathbb{s}_{1}$ (where $A_{i}$ and $a_{j}$ are defined over $\mathbb{k}(\mathbb{s})$). Interesting examples of such families arise as limits of periodic families when some of the parameters $s_{j}$ are replaced by $s_{j}/t$ and $t$ is sent to 00, see Section 4 below.
The following generalization of Theorem 3.3 to the case of mixed-periodic families is obtained by a straighforward generalization (or, more precisely, degeneration) of its proof. Let
$\mathbb{B}(\nabla(\mathbb{s})):=\sum_{j=1}^{n_{1}}(s_{j}-s_{j}^{p})B_{j}^{(1)}-\sum_{j=n_{1}+1}^{n}s_{j}^{p}B_{j}^{(1)}.$
Theorem 3.7.
If $\nabla(\mathbb{s})$ is a mixed-periodic pencil (periodic in the first $n_{1}$ variables $s_{j}$ and infinitesimally periodic in the next $n_{2}$ variables $s_{j}$) then the pencils $C(\nabla(\mathbb{s}))$ and $\mathbb{B}(\nabla(\mathbb{s}))$ of commuting endomorphisms are isospectral for all $\mathbb{s}\in\mathbb{k}^{n}$. In particular, if $s_{j}\in\mathbb{F}_{p}$ for all $1\leq j\leq n_{1}$ and $s_{j}=0$ for $n_{1}+1\leq j\leq n$ then the endomorphisms $C_{i}(\nabla(\mathbb{s}))$ are nilpotent. Moreover, if for some $\mathbb{s}$, $\mathbb{B}(\nabla(\mathbb{s}))$ has simple spectrum, then $C(\nabla(\mathbb{s}))$ and $\mathbb{B}(\nabla(\mathbb{s}))$ are conjugate.
Proof.
The proof is completely parallel to the proof of Theorem 3.3; the only significant change is that besides the fact that a polynomial $b(s)$ in characteristic $p$ such that $b(s+1)=b(s)$ is a polynomial of $s^{p}-s$, we should also use the fact that if $b^{\prime}(s)=0$ then $b$ is a polynomial of $s^{p}$. ∎
3.4. Global nilpotence and regularity of periodic pencils
Definition 3.8.
(N. Katz, [K1, K2]) A flat connection $\nabla$ defined over $\overline{\mathbb{Q}}$ is said to be globally nilpotent if the $p$-curvature of its reduction to characteristic $p$ is nilpotent for almost all $p$.
Corollary 3.9.
If normal-$\nabla$ is a periodic pencil defined over $\overline{\mathbb{Q}}$ then for any $\mathbb{s}\in\mathbb{Q}^{n}$, the connection $\nabla(\mathbb{s})$ is globally nilpotent. In particular, this is so for the pencils in Theorem 3.11.
Proof.
This follows immediately from Theorem 3.3. ∎
Let ${\rm char}(\mathbb{k})=0$. Recall that a family of flat connections $\nabla$ over $\mathbb{k}$ is quasi-motivic if it is periodic and has regular singularities ([EV], Definition 4.3).
Corollary 3.10.
Any periodic pencil $\nabla(\mathbb{s})$ over $\mathbb{k}$ has regular singularities, and also has rational residues for $\mathbb{s}\in\mathbb{Q}^{n}$ (hence quasiunipotent monodromies for such $\mathbb{s}$ if $\mathbb{k}=\mathbb{C}$). In particular, the notions of a periodic and a quasi-motivic family in the special case of pencils are equivalent.
Proof.
Without loss of generality we may assume that $\nabla$ is defined over $\overline{\mathbb{Q}}$. By a theorem of N. Katz conjectured by Grothendieck ([K2], Section 13), any globally nilpotent linear differential equation defined over $\overline{\mathbb{Q}}$ has regular singularities. Thus by Corollary 3.9 $\nabla(\mathbb{s})$ has regular singularities for $\mathbb{s}\in\mathbb{Q}^{n}$. Pick a smooth compactification $\overline{X}$ of $X$ with normal crossing divisor $D$ at $\infty$. Since the singular locus of $\nabla(\mathbb{s})$ on $\overline{X}$ is $D$ for all $\mathbb{s}\in\mathbb{k}^{n}$ (i.e., they are not moving as $\mathbb{s}$ varies), it follows that the singularities of $\nabla(\mathbb{s})$ are regular for all $\mathbb{s}\in\mathbb{k}^{n}$. ∎
3.5. Applications of the main theorem
Theorem 3.11.
Let $\nabla(\mathbb{s})$ be any of the pencils in [EV], Section 5 (KZ, Dunkl, and Casimir connections), or the equivariant quantum connection for a conical symplectic resolution with finitely many torus fixed points ([EV], Subsection 3.5). Then for almost all $p$, the $p$-curvature of a reduction $d-\sum_{j=1}^{n}s_{j}B_{j}$ of $\nabla(\mathbb{s})$ at $p$ is isospectral to $\sum_{j=1}^{n}(s_{j}-s_{j}^{p})B_{j}^{(1)}$. In particular, if $s_{j}\in\mathbb{F}_{p}$ for all $j$ then this $p$-curvature is nilpotent.
Proof.
It is shown in [EV] that these pencils are periodic, so the result immediately follows from Theorem 3.3. ∎
Remark 3.12.
More specifically, we see that Theorem 3.11 holds for those reductions of $\nabla$ to characteristic $p$ which come with reductions of the shift operators $A_{j}$. So if $A_{j}$ are known explicitly (or at least defined over an explicit finitely generated ring, as in the case of quantum connections where $A_{j}$ are constructed geometrically) then we can produce a list of forbidden primes away from which Theorem 3.11 holds. However, in cases when the existence of $A_{j}$ is proved non-constructively (e.g., using that the pencil is regular and has periodic monodromy, as in the examples of [EV], Section 5) we cannot specify a lower bound for $p$ sufficient for Theorem 3.11.
Remark 3.13.
Theorem 3.11 allows one to compute (albeit somewhat implicitly) the eigenvalues of the $p$-curvature in many of the above examples. For instance, for the KZ connection $d-\hbar\sum_{i\neq j}\frac{\Omega_{ij}}{x_{i}-x_{j}}dx_{i}$ on the tensor product of finite dimensional representations of a Lie algebra, Theorem 3.11 says that the $p$-curvature is isospectral to the Gaudin operators renormalized by the factor pPlanck-constant-over-2-pisuperscriptPlanck-constant-over-2-$\hbar-\hbar^{p}$, so its eigenvalues can be found using the Bethe Ansatz method.
3.6. Motivic families
Recall that a family $\nabla$ of flat connections on $X$ is said to be motivic if there exists an irreducible variety $Y$, a smooth morphism $\pi:Y\to X$, and non-vanishing regular functions $\Phi_{1},...,\Phi_{n}$ on $Y$ such that on some dense open set $X^{\circ}\subset X$, for Zariski generic $\mathbb{s}$ the connection $\nabla(\mathbb{s})$ is isomorphic to the Gauss-Manin connection on $H^{i}(\pi^{-1}(\mathbb{x}),\mathcal{L}(\Phi_{1},..,\Phi_{n},\mathbb{s}))$ for some $i$, where $\mathcal{L}(\Phi_{1},..,\Phi_{n},\mathbb{s})$ is the local system generated by the multivalued function $\prod_{j=1}^{n}\Phi_{j}^{s_{j}}$ ([EV], Definition 4.14). By [EV], Proposition 4.15, this is a subclass of quasi-motivic families.
Also recall Katz’s theorem:
Theorem 3.14.
([K1]) The Gauss-Manin connection on the cohomology of fibers of a smooth morphism defined over $\overline{\mathbb{Q}}$ is globally nilpotent.
Katz’s theorem implies that Corollary 3.9 holds for motivic families (not necessarily pencils):
Proposition 3.15.
If normal-$\nabla$ is a motivic family then $\nabla(\mathbb{s})$ is globally nilpotent for any $\mathbb{s}\in\mathbb{Q}^{n}$.
Proof.
For $\mathbb{s}\in\mathbb{Q}^{n}$ the function $\prod_{j=1}^{n}\Phi_{j}^{s_{j}}$ can be viewed as a regular function on a finite cover $\widetilde{Y}$ of $Y$ whose degree is the common denominator of $s_{j}$, so the corresponding connection is the Gauss-Manin connection on the cohomology of the fibers of the map $\widetilde{\pi}:\widetilde{Y}\to X$. Thus the result follows by Katz’s theorem. ∎
Remark 3.16.
There is another proof of Theorem 3.3 for 1-parameter motivic pencils $\nabla(s)$ attached to a multivalued function $\Phi^{s}$ on $Y$, provided that the “master function” $\Phi$ has $\dim H^{\dim Y-\dim X}(\pi^{-1}(\mathbb{x}),\mathcal{L}(\Phi,s))$ nondegenerate critical points on fibers of $\pi:Y\to X$ (this happens, for instance, for the KZ connection for $\mathfrak{sl}_{2}$). This proof, proposed by the second author and V. Vologodsky ([VV]), is based on the Hodge-theoretic methods in the spirit of [K1], showing that the $p$-curvature of $\nabla(s)$ reduced to characteristic $p$ “localizes” in a suitable sense to the critical points of $\Phi$. So this proof has the advantage of not using the linearity of $\nabla$ in $s$, and therefore extends to the more general case of motivic families (not just pencils) of flat connections. We hope that this method of proof can also be generalized to multiparameter families and to cases when critical points of $\Phi$ are degenerate, or even non-isolated.
3.7. The André-Bombieri-Dwork conjecture
The André-Bombieri-Dwork conjecture [A] states that if a first order system of linear differential equations on $\mathbb{P}^{1}$ defined over ¯$\overline{\mathbb{Q}}(z)$ is globally nilpotent, then it is geometric, i.e., its solutions admit an integral representation. While we can’t offer any definitive ideas towards a proof or disproof of this conjecture, our results provide numerous examples of concrete globally nilpotent systems for which a geometric construction is not known (even though in some cases it has been sought for a long time).
For instance, such examples include rational and trigonometric Dunkl connections for exceptional complex reflection groups ([EV], Subsection 5.3) and elliptic KZ connections ([EV], Subsection 5.1.7), for rational values of the parameters $s_{j}$. By Corollary 3.9, they are all globally nilpotent, and can be made into connections on $\mathbb{P}^{1}$ by restricting to any line (or, more generally, rational curve) on the base.
For instance, there is no known geometric construction of the (rational or trigonometric) Dunkl connection $\nabla(c)$ for exceptional groups $W=E_{6},E_{7},E_{8}$. In other words, there is no known integral representation for solutions of this system (Heckman-Opdam hypergeometric functions for $W$), even though considerable efforts have been made to find one. See e.g. the discussion at the beginning of [CHL].
4. Irregular pencils
We would now like to apply the above results to computing the spectrum of the $p$-curvature for irregular connections. Many interesting examples of such connections arise as confluent limits $\nabla=\lim_{t\to 0}\nabla_{t}$ of 1-parameter families of regular connections for which the spectrum of the $p$-curvature has already been computed. In such cases, we may obtain the spectrum of the $p$-curvature of $\nabla$ from that of $\nabla_{t}$ by a limiting procedure, using that this spectrum depends on $t$ algebraically. Let us list some examples which can be handled in this way.
4.1. Irregular KZ connections
The KZ connection of [EV], Subsection 5.1 admits an irregular deformation considered in [FMTV]. Namely, as in [EV], Section 5.1, let $\mathfrak{g}$ be a finite dimensional simple Lie algebra over $\mathbb{C}$, $\Omega\in S^{2}\mathfrak{g}$ the Casimir tensor, $\mathfrak{h}\subset\mathfrak{g}$ a Cartan subalgebra, $h\in\mathfrak{h}$, and consider the connection
(4.1) $\nabla(\hbar,h):=d-\sum_{i=1}^{r}\left(h^{(i)}+\hbar\left(\sum_{j\neq i}\frac{\Omega^{ij}}{x_{i}-x_{j}}\right)\right)dx_{i}$
on the trivial bundle on $\mathbb{C}^{r}\setminus{\rm diagonals}$ with fiber being the weight space $(V_{1}\otimes...\otimes V_{r})[\mu]$, where $V_{i}$ are finite dimensional representations of $\mathfrak{g}$. This connection is obtained as a limiting case of the trigonometric KZ connection of [EV], Subsection 5.1.6 (for the usual quasitriangular structure $\mathbb{r}$ on $\mathfrak{g}$) by replacing $x_{i}$ with $1+tx_{i}$, setting $s:=h/t$, and sending $t$ to 00.
We thus obtain
Proposition 4.1.
For almost all $p$, the $p$-curvature operators $C_{i}(\hbar,h)(\mathbb{x})$ of the reduction of $\nabla(\hbar,h)$ at $p$ are isospectral to
$C_{i}^{*}(\hbar,h)(\mathbb{x}):=(-h^{p})^{(i)}+(\hbar-\hbar^{p})\sum_{j\neq i}\frac{\Omega^{ij}}{x_{i}^{p}-x_{j}^{p}}.$
Remark 4.2.
Note that this is slightly different from the expression in Theorem 3.3: in the first term instead of $h-h^{p}$ we have just $-h^{p}$, as in Theorem 3.7. This is because we have set $s=h/t$ and sent $t$ to 00. In fact, one can show that the limiting pencil $\nabla(\hbar,h)$ is mixed-periodic: it is periodic in $\hbar$ and infinitesimally periodic in $h$. Namely, the connection with respect to $h$ commuting with $\nabla(\hbar,h)$ is exactly the dynamical (or Casimir) connection of [FMTV], which is bispectrally dual to the irregular KZ connection.
4.2. Irregular Casimir connections
The Casimir connection with values $(V_{1}\otimes...\otimes V_{n})[\mu]$ ([EV], 5.3.1) admits an irregular deformation, which was introduced in [FMTV] and is (as just noted in Remark 4.2) bispectrally dual to the irregular KZ connection. To define it, we need to equip $V_{1}\otimes...\otimes V_{n}$ with the structure of a representation of the Lie algebra $\mathfrak{g}[z]$, by considering the tensor product of evaluation representations $V_{1}(x_{1})\otimes...\otimes V_{n}(x_{n})$, where $\mathbb{x}=(x_{1},...,x_{n})\in\mathbb{C}^{n}$. Let $\mathfrak{h}_{\rm reg}$ be the set of regular elements of $\mathfrak{h}$, $\alpha_{i},\omega_{i}^{\vee}$ be the simple roots and fundamental coweights of $\mathfrak{g}$, $R_{+}$ be the set of positive roots of $\mathfrak{g}$, $e_{\alpha},e_{-\alpha}\in\mathfrak{g}$ be the root elements corresponding to $\alpha\in R_{+}$. Then the irregular Casimir connection is the connection on the trivial bundle over ${\mathfrak{h}}_{\rm reg}$ with fiber $V:=(V_{1}(x_{1})\otimes...\otimes V_{n}(x_{n}))[\mu]$ given by
$\nabla(\hbar,\mathbb{z})=d-\hbar\sum_{\alpha\in R_{+}}\frac{e_{\alpha}e_{-\alpha}+e_{-\alpha}e_{\alpha}}{2}\frac{d\alpha}{\alpha}+\sum_{i=1}^{{\rm rank}\mathfrak{g}}(\omega_{i}^{\vee}\otimes z)d\alpha_{i},$
where $\omega_{i}^{\vee}\otimes z\in\mathfrak{g}[z]$ acts in $V$. This connection is a composition factor of the limit of the trigonometric Casimir connection ([EV], Subsection 5.3.2) under rescaling $h\mapsto th$ for $h\in\mathfrak{h}$, $\mathbb{x}\mapsto\hbar\mathbb{x}/t$, and sending $t$ to 00, which corresponds to degeneration of the Yangian $Y(\mathfrak{g})$ to $U(\mathfrak{g}[z])$ (it is a composition factor rather than the whole limit because the representations $V_{j}$ may not lift to $Y(\mathfrak{g})$).
Let $C_{i}(\hbar,\mathbb{x})$ be the $p$-curvature operator of the reduction of $\nabla(\hbar,\mathbb{x})$ at $p$ corresponding to the vector $\omega_{i}^{\vee}\in\mathfrak{h}$. Then we get
Proposition 4.3.
For almost all $p$, the operators $C_{i}(\hbar,\mathbb{x})(h)$ are isospectral to
$C_{i}^{*}(\hbar,\mathbb{x})(h):=(\hbar-\hbar^{p})\sum_{\alpha\in R_{+}}\frac{\alpha(\omega_{i}^{\vee})}{\alpha(h)^{p}}\frac{e_{\alpha}e_{-\alpha}+e_{-\alpha}e_{\alpha}}{2}+\sum_{j=1}^{n}x_{j}^{p}(\omega_{i}^{\vee})^{(j)}.$
Remark 4.4.
Similarly to the irregular KZ connection, one can show that the irregular Casimir connection is periodic in $\hbar$ and infinitesimally periodic in $x_{j}$, with the commuting (bispectral dual) connection with respect to $\mathbb{x}$ being the irregular KZ connection.
4.3. Irregular Dunkl connections
Let $W$ be a finite Coxeter group with set of reflections $S$ and reflection representation $\mathfrak{h}$ of dimension $r$. Let $\mathbb{c}:S\to\mathbb{C}$ be a $W$-invariant function. The Dunkl connection ([EV], Subsection 5.2) with $V=\mathbb{C}W$ admits an irregular deformation. Namely, for $\lambda\in\mathfrak{h}^{*}$ let $V_{\lambda}^{0}$ be the representation of $\mathbb{C}W\ltimes S\mathfrak{h}$ on $\mathbb{C}W$ where $W$ acts by left multiplication and $h\circ w=\lambda(w^{-1}h)w$, $w\in W$, $h\in\mathfrak{h}$. The irregular Dunkl connection is then the connection over the regular locus $\mathfrak{h}_{\rm reg}\subset\mathfrak{h}$ with fiber $V_{\lambda}^{0}$ given by
(4.2) $\nabla(\mathbb{c},\lambda)=d-\sum_{i=1}^{r}u_{i}du_{i}^{*}-\sum_{w\in S}\mathbb{c}(w)\frac{d\alpha_{w}}{\alpha_{w}}w,$
where $\alpha_{w}$ is a root corresponding to a reflection $w\in S$, $\{u_{i}\}$ is a basis of $\mathfrak{h}$, $u_{i}^{*}$ the dual basis of $\mathfrak{h}^{*}$, and $u_{i}$ acts in $V_{\lambda}^{0}$. If $W$ is a Weyl group then this is a limiting case of the trigonometric Dunkl connection $\nabla(\mathbb{c})|_{V_{\lambda/t}}$ ([EV], Subsection 5.2.2) by zooming in with $h\mapsto th$, $h\in\mathfrak{h}$, and sending $t$ to 00.
We thus obtain
Proposition 4.5.
If $W$ is the Weyl group of a root system $R$ then for almost all $p$, the $p$-curvature operators $C_{i}(\mathbb{c},\lambda)(h)$ of the reduction of $\nabla(\mathbb{c},\lambda)$ at $p$ corresponding to the fundamental coweights $\omega_{i}^{\vee}\in\mathfrak{h}$ are isospectral to
$C_{i}^{*}(\mathbb{c},\lambda)(h):=\left(-\omega_{i}^{\vee}+\sum_{w\in S}\mathbb{(}\mathbb{c}-\mathbb{c}^{p})(w)\frac{\alpha_{w}(\omega_{i}^{\vee})}{\alpha_{w}^{p}(h)}w\right)\biggl{|}_{V_{\lambda^{p}}^{0}}.$
Remark 4.6.
As before, one can show that the irregular Dunkl connection is periodic in $\mathbb{c}$ and infinitesimally periodic in $\lambda$, and is bispectrally self-dual: the bispectrally dual connection is the irregular Dunkl connection with $\lambda$ and $h$ swapped.
Remark 4.7.
The irregular Dunkl connection (4.2) is a rare case of a nontrivial mixed-periodic pencil for which all the (infinitesimal) shift operators $A_{j},a_{k}$ can be computed fairly explicitly. Namely, the irregular Dunkl connection can be described as the flat connection attached to the eigenvalue problem for the rational Calogero-Moser operators:
[…]
In this realization the shift operators $A_{j}$ are the Dunkl-Opdam shift operators $S_{j}(\mathbb{c})$ ([DO], Subsection 3.5), which are explicit differential operators with rational coefficients which satisfy the identities
$L_{k}(\mathbb{c}+\mathbb{e}_{j})\circ S_{j}(\mathbb{c})=S_{j}(\mathbb{c})\circ L_{k}(\mathbb{c})$
and hence map joint eigenfunctions of $L_{k}(\mathbb{c})$ to joint eigenfunctions of $L_{k}(\mathbb{c}+\mathbb{e}_{j})$, thus defining an isomorphism of the corresponding $D$-modules. This can be used to show that an analog of Proposition 4.5 holds for all finite complex reflection groups (not just Weyl groups). One just needs to remember that reflection representations of complex reflection groups may be defined not over $\mathbb{Q}$ but over some number field $K$. This needs to be taken into account when computing Frobenius twists.
The same holds for the trigonometric Dunkl connection ([EV], Subsection 5.2.2) using its realization as the eigenvalue problem for trigonometric Calogero-Moser operators. Namely, the shift operators in this case are Opdam’s shift operators ([O]).
4.4. Toda connections of finite type
Toda connections are a special confluent limit of Dunkl-Cherednik connections (see e.g. [BMO], Section 7). To explain this limit, for simplicity assume that the root system $R$ is irreducible and simply laced. Recall from [EV], Subsection 5.2.2 that in this case we have the degenerate affine Hecke algebra $\mathcal{H}_{c}$, $c\in\mathbb{C}$, generated by the Weyl group $W$ and $\mathfrak{h}$ with defining commutation relations
$s_{i}h-s_{i}(h)s_{i}=c\alpha_{i}(h),\ h\in\mathfrak{h}$
for simple reflections $s_{i}$ and simple roots $\alpha_{i}$, and for $\lambda\in\mathfrak{h}^{*}$ we define the induced representation $V_{c,\lambda}:=\mathcal{H}_{c}\otimes_{\rm Sym{\mathfrak{h}}}\mathbb{C}_{\lambda}$. Consider the torus $H=\mathfrak{h}/2\pi iP$ where $P$ is the (co)weight lattice. Recall that the Dunkl-Cherednik connection is the connection on the trivial bundle over the regular locus $H_{\rm reg}$ of $H$ with fiber $V_{c,\lambda}$ given by
$\nabla(c,\lambda)=d-\sum_{i=1}^{r}\omega_{i}d\alpha_{i}-c\sum_{\alpha\in R_{+}}\frac{e^{\alpha}d\alpha}{1-e^{\alpha}}(s_{\alpha}-1),$
where $\omega_{i}$ are the fundamental (co)weights.
In the Toda limit, the parameter $\lambda$ is kept fixed, while the parameter $c$ is sent to infinity. But in order for the limit to exist, we must also make a shift on $H$ and send it to infinity in a coordinated way with $c$, so the Weyl group symmetry is lost. To explain how exactly this works, introduce renormalized simple reflections $\overline{s}_{i}:=c^{-1}s_{i}$. They satisfy the defining relations $\overline{s}_{i}^{2}=c^{-2},\ \overline{s}_{i}h-s_{i}(h)\overline{s}_{i}=\alpha_{i}(h),\ h\in\mathfrak{h}$, and the braid relations. Thus in the limit $c\to\infty$ these elements satisfy the relations
$\overline{s}_{i}^{2}=0,\ \overline{s}_{i}h-s_{i}(h)\overline{s}_{i}=\alpha_{i}(h),\ h\in\mathfrak{h}$
and the braid relations. These are defining relations of the nil-Hecke algebra $\mathcal{H}_{\infty}$. The representation $V_{c,\lambda}$ of $\mathcal{H}_{c}$ degenerates into the representation $V_{\infty,\lambda}:=\mathcal{H}_{\infty}\otimes_{\rm Sym{\mathfrak{h}}}\mathbb{C}_{\lambda}$ of $\mathcal{H}_{\infty}$.
Now we are ready to take the Toda limit. For this purpose, introduce renormalized reflections $\overline{s}_{\alpha}=c^{-\ell(s_{\alpha})}s_{\alpha}$, $\alpha\in R_{+}$, where $\ell$ denotes the length (these elements have nonzero limits in $\mathcal{H}_{\infty}$). In terms of these elements, we can write the Dunkl-Cherednik connection as
$\nabla(c,\lambda)=d-\sum_{i=1}^{r}\omega_{i}d\alpha_{i}-c\sum_{\alpha\in R_{+}}\frac{e^{\alpha}d\alpha}{1-e^{\alpha}}(c^{\ell(s_{\alpha})}\overline{s_{\alpha}}-1),$
We see that this has no limit as $c\to\infty$ unless we make a shift along $H$. So let us make the shift $g\mapsto gc^{-2\rho}$, $g\in H$. After this we have
$\nabla(c,\lambda)=d-\sum_{i=1}^{r}\omega_{i}d\alpha_{i}-c\sum_{\alpha\in R_{+}}\frac{c^{-2|\alpha|}e^{\alpha}d\alpha}{1-c^{-2|\alpha|}e^{\alpha}}(c^{\ell(s_{\alpha})}\overline{s_{\alpha}}-1),$
where $|\alpha|$ is the height of $\alpha$. But it is known that $\ell(s_{\alpha})=2|\alpha|-1$ (see [BMO] and references therein), so we get
$\nabla(c,\lambda)=d-\sum_{i=1}^{r}\omega_{i}d\alpha_{i}-\sum_{\alpha\in R_{+}}\frac{e^{\alpha}d\alpha}{1-c^{-2|\alpha|}e^{\alpha}}(\overline{s_{\alpha}}-c^{-\ell(s_{\alpha})}).$
Now this has a finite limit as $c\to\infty$ given by
$\nabla(\infty,\lambda)=d-\sum_{i=1}^{r}\omega_{i}d\alpha_{i}-\sum_{\alpha\in R_{+}}e^{\alpha}\overline{s_{\alpha}}d\alpha.$
This limit is the Toda connection of the root system $R$. Its base is now the whole torus $H$ (the singularities disappeared in the limit), but it has irregular singularities at infinity.
Since $\nabla(c,\lambda)$ is periodic in $\lambda$, so is $\nabla(\infty,\lambda)$; in fact, as was shown by A. Givental, this connection arises as the equivariant quantum connection for the flag variety $G/B$ of the simple complex Lie group $G$ corresponding to $R$, so the shift operators for $\nabla(\infty,\lambda)$ can be constructed geometrically. Moreover, they can be computed explicitly either by taking the limit of Opdam shift operators, or from geometry. In fact, this was done in type $A$ in the more general case of partial flag varieties in [TV].
However, in the terminology of [EV], $\nabla(\infty,\lambda)$ is not a pencil but rather an affine pencil, in the sense that its dependence on $\lambda$ is linear, but inhomogeneous (owing to the last summand). This means that we cannot compute the spectrum of the $p$-curvature of $\nabla(\infty,\lambda)$ using Theorem 3.3.
But we can do this computation using the above limiting procedure and the knowledge of the spectrum of the $p$-curvature of the Dunkl-Cherednik connection. Namely, by Theorem 3.3, the $p$-curvature operators $C_{i}(c,\lambda)$ corresponding to $\omega_{i}$ for the reduction of $\nabla(c,\lambda)$ at $p$ are isospectral to
$C_{i}^{*}(c,\lambda)=-(\omega_{i}+(c^{p}-c)\sum_{\alpha\in R_{+}}(\alpha,\omega_{i})\tfrac{c^{-2p\rho}e^{p\alpha}}{1-c^{-2p\rho}e^{p\alpha}}(s_{\alpha}-1))\big{|}_{V_{c^{p}-c,\lambda^{p}-\lambda}}.$
So rewriting this in terms of $\overline{s}_{i}=\frac{s_{i}}{c^{p}-c}$ (as $c$ has been replaced by $c^{p}-c$) and taking the limit $c\to\infty$, we obtain the following statement.
Proposition 4.8.
The $p$-curvature operators $C_{i}(\infty,\lambda)(g)$ corresponding to $\omega_{i}$ for the reduction of $\nabla(\infty,\lambda)$ at $p$ are isospectral to
$C_{i}^{*}(\infty,\lambda)(g)=-(\omega_{i}+\sum_{\alpha\in R_{+}}(\alpha,\omega_{i})\alpha(g)^{p}\overline{s}_{\alpha})\big{|}_{V_{\infty,\lambda^{p}-\lambda}},\ g\in H.$